OpenStudy (anonymous):

What is the equation of the hyperbola with vertices at (0,-4) and (0,4) and foci at (0,-6) and (0,6)? a. y^2/16 - x^2/20 =1 b. x^2/16 - y^2/20 =1 c. x^2/16 - y^2/36 =1 d. x^2 - y^2/52 =1

OpenStudy (mertsj):

a^2=16, c^2=36 a^2+b^2=c^2 If the vertices are (0,4) and (0,-4), what is the center?

OpenStudy (anonymous):

2? I have no idea

OpenStudy (mertsj):

do you know what a hyperbola is?

OpenStudy (anonymous):

yes sort of

OpenStudy (mertsj):

Do you know what the vertex of a hyperbola IS?

OpenStudy (anonymous):

no

OpenStudy (mertsj):